Currently, the one year spot rate is 0.50% per year and the two year spot rate is 1.00% per year. What is the expected one-year spot rate starting one year from today under the Pure Expectations Theory?
Answer #1

We know that expected one year spot rate one year from today is given as equal to=(1+2 year spot rate)^2/(1+1 year spot rate)-1=1.01^2/1.005-1=1.50%